OpenStudy (anonymous):

Solve -1/2x≥-7/4 (simplify your answer; type an inequality symbol then type an integer or fraction)

OpenStudy (anonymous):

\[-1/2\ge-7/4\] \[-0.5x \ge-1.75\] \[x \ge-1.75/-0.5\] \[x \ge3.5\]

OpenStudy (anonymous):

Thank you, I am showing that is incorrect though, :(

OpenStudy (anonymous):

show me than